### v1

@@ 2.2. Computation of the spectral flow: reduction to the Jacobi case
[…]
@@
We will take advantage of the robustness of the spectral flow and reduce the problem to simpler models. For instance, we will first prove the results for $1$–periodic Jacobi operators (for which $h(n)=0$ for $|n|\geq 2$), and then extend it to all $1$–periodic Hamiltonians by density. The proof that the spectral flow equals $-{\mathcal{N}}(E)$ for Jacobi operator is postponed to the next Section. In this section, we explain how to generalize the result for Jacobi operators to any periodic Hamiltonian.
@@ A Jacobi operator is an operator $H$ acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ of the form
We will establish a version of Theorem 1.1 in the special of Jacobi operators.
@@
If $\widetilde{H}$ denotes the corresponding operator on $\ell^{2}(\mathbb{Z},\mathbb{C}^{\ell N})$, then $\sigma(H)=\sigma(\widetilde{H})$, and each Bloch band of $H$ becomes a set of $\ell$ bands for $\widetilde{H}$. In particular, ${\mathcal{N}}_{\widetilde{H}}(E)=\ell{\mathcal{N}}_{H}(E)$.
@@
$\widetilde{w}(x):=\begin{pmatrix}w(\ell x)&0&\cdots&0\\ 0&w(\ell x+1)&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&w(\ell x+\ell-1)\end{pmatrix}$
@@
${\mathrm{Sf}}\left(\widetilde{H}^{\sharp}(t),[0,1],E\right)={\mathrm{Sf}}\left(H^{\sharp}(t),[0,\ell],E\right)=\ell\,{\mathrm{Sf}}\left(H^{\sharp}(t),[0,1],E\right).$
@@
${\mathrm{Sf}}\left(\widetilde{H}^{\sharp}(t),[0,1],E\right)=-{\mathcal{N}}_{\widetilde{H}}(E)=-\ell{\mathcal{N}}_{H}(E),$
@@
This kernel has already been introduced in (7). Recall that $\|H-H_{\ell}\|_{\rm op}\to 0$ by Young’s inequality. Note also that $\|H^{\sharp}(t)-H_{\ell}^{\sharp}(t)\|_{\rm op}=\|H-H_{\ell}\|_{\rm op}$ is independent of $t$, hence goes to 00 uniformly in $t\in\mathbb{R}$. In particular, the spectrum of $H_{\ell}^{\sharp}(t)$ approaches uniformly the spectrum of $H^{\sharp}(t)$.
@@
${\mathrm{Sf}}\left(H^{\sharp}(t),[0,1],E\right)=\lim_{\ell\to\infty}{\mathrm{Sf}}\left(H^{\sharp}_{\ell}(t),[0,1],E\right)=-{\mathcal{N}}(E),$
@@
The main argument is that since $t\mapsto H^{\sharp}(t)$ is $\nu$–Lipschitz, so are all the eigenvalues branches. Let $E\in\mathbb{R}\setminus\sigma_{\rm{bulk}}$ with ${\mathcal{N}}(E)>0$. Since there is a nonzero spectral flow, $H^{\sharp}(t)$ has eigenvalues for at least some values of $t$. By shifting $t$ if necessary, we may assume $t=0$. We label this eigenvalue $\lambda_{0}(0)$ and use this as a starting point to label the eigenvalue curves $\lambda_{k}(t)$ for $t\in[0,1]$ such that, for all $t\in[0,1]$
@@
and with the convention that $\lambda_{k}(t)$ equals the edge of the essential spectrum if there exists no $k$-th eigenvalue at this value of $t$. Since the spectrum is periodic, we have $\lambda_{k}(1)=\lambda_{j}(0)$ for some $j\in\mathbb{Z}$. The spectral flow equals $-{\mathcal{N}}(E)$, so there are exactly ${\mathcal{N}}(E)$ eigenvalues that crosses $\lambda_{k}(1)$ upwards. We deduce that we actually have $\lambda_{k}(1)=\lambda_{k+{\mathcal{N}}(E)}(0)$ for all values of $k\in\mathbb{Z}$ such that $\lambda_{k}(0)$ is an eigenvalue of $H^{\sharp}(0)$. Then, the Lipschitz continuity of $t\mapsto\lambda_{k}(t)$ gives
$\lambda_{k+{\mathcal{N}}(E)}(0)-\lambda_{k}(0)=\lambda_{k}(1)+\lambda_{k}(0)\leq\nu.$
Hence, there are at least ${\mathcal{N}}(E)$ eigenvalues for $H^{\sharp}(0)$ in each interval of length $\nu$ in the gap. If $\nu$ is larger than the width of the gap, the same argument shows that there is edge spectrum for any $t_{0}\in[0,1]$, and we can repeat the counting argument for this $t_{0}$.
@@ 2.4.2. Numerical simulations for the SSH chain
which corresponds to the evaluation of $w_{1}$ at the location of each of the two atoms in the first unit cell. Note that $w$ satisfy the assumption of a soft-wall, and is $\nu$–Lipschitz.
@@
Then, for each eigenpair $(\lambda,u)$ with $\|u\|_{\ell^{2}_{L}}=1$, we check if the mass of $u$ is localized near the origin, or, more specifically, not localized at the boundary of our simulation box. This procedure discards the spurious edge modes due to the truncation. We detect that $u$ is an true edge mode whenever $\|u{\mathds{1}}(-L/2\leq n\leq L/2)\|_{\ell^{2}_{L}}\geq 3/4$.
@@
Figure 3. Numerics for the spectrum of $t\mapsto H^{\sharp}(t)$ in the SSH model, for different values of $\nu$. From left to right, $\nu=0.5$, $\nu=1$, $\nu=5$ and $\nu=10$. The essential spectrum is displayed in grey, and the edge spectrum in red. For $E$ in the gap between the $2$ bands, ${\mathcal{N}}(E)=1$ and the spectral flow is $-1$, above the bands, ${\mathcal{N}}(E)=2$ and the spectral flow equals $-2$.
@@ 3. Computation of spectral flows in the Jacobi case
Before we start, we emphasize that several techniques are available to study Jacobi operators, such as the ones involving transfer matrices. Recall indeed that if $H$ is Jacobi, then the equation $H\Psi=E\Psi$ implies that
@@
so that we can study the properties of $\Psi$ by studying the transfer matrix $T(E)$. However, this matrix is not always well-defined (in the previous formulation, we need $a$ to be invertible for instance). Actually, we were not able to recover our results using the formalism of transfer matrices. Our proof rather relies on a direct study of the full operator, and a variational argument.
@@ 3.1. Reduction to the cut case
[…]
@@
The difference of potentials $w-w_{\Sigma}$ is continuous, vanishes for $x<x_{\Sigma}$, and goes to 00 as $x\to+\infty$. In particular, for all $t\in\mathbb{R}$, $(W-W_{\Sigma})(t)$ is compact, and the map $t\mapsto(W-W_{\Sigma})(t)$ is continuous and $1$–quasi–periodic.
@@
$H^{\Sigma}(t)-H^{\sharp}(t)=W(t)-W_{\Sigma}(t),$
which is compact for all $t\in\mathbb{R}$. In addition, both operators $H^{\Sigma}$ and $H^{\sharp}$ are $1$–quasi–periodic. By robustness of the spectral flow with respect to compact operators, see Lemma A.11, we deduce that
@@
Note that $K(1)=\tau_{1}K(0)\tau_{1}^{*}$, so the map $t\mapsto K(t)$ can be extended by $1$–quasi–periodicity, and this extension is continuous. For simplicity, in what follows, we will restrict the study to $t\in[0,1]$.
@@
$H_{\rm{cut}}^{\Sigma}(t):=H^{\Sigma}(t)-K(t),\quad\text{so that}\quad H^{\Sigma}(t)=H^{\Sigma}_{\rm{cut}}(t)+K(t)$
Since $t\mapsto H^{\Sigma}(t)$ and $t\mapsto K(t)$ are continuous and $1$–quasi–periodic, so is $t\mapsto H_{\rm{cut}}^{\Sigma}(t)$. So, using again the robustness of the spectral flow with respect to compact, quasi-periodic operators, see Lemma A.11, we find that
@@
These left and right maps are not periodic nor $1$-quasi-periodic in $t$. Still, one can define the spectral flow as the net number of eigenvalues crossing the energy $E$ downwards. Since these operators are no longer $1$–quasi–periodic, a priori this spectral flow could depend on the value of $E$ in the essential gap (see Appendix A). In any case, since $H_{\rm{cut}}^{\Sigma}(t)$ is block diagonal, we have
@@
$H^{\Sigma}_{R}(0)=\left(\begin{array}[]{| c c c c }\hline\cr b&a&&\\ a&b&a&\\ &a&b&\ddots\\ &&\ddots&\ddots\end{array}\right),\qquad\text{and}\qquad H^{\Sigma}_{R}(1)=\left(\begin{array}[]{| c c c c }\hline\cr\Sigma&0&&\\ 0&b&a&\\ &a&b&\ddots\\ &&\ddots&\ddots\end{array}\right).$
As noted before, the half-line operators are no longer $1$–quasi–periodic: the presence of the boundary at 00 makes the model at $t=1$ not unitarily equivalent to the one at $t=0$. The key observation is that the relevant part of the spectrum is still periodic, since $H^{\Sigma}_{R}(1)$ is a direct sum of $\Sigma$ with (a shifted version of) $H^{\Sigma}_{R}(0)$. We will study such families of operators in the next section.

### v3

@@ 2.2. Computation of the spectral flow: reduction to the Jacobi case
$\forall E>0,\qquad{\mathrm{Sf}}(W(t),E,[0,1])=-N,\qquad\text{while}\qquad\forall E<0,\qquad{\mathrm{Sf}}(W(t),E,[0,1])=0.$
@@
We will take advantage of the robustness of the spectral flow and reduce the problem to simpler models. For instance, we will first prove the results for $1$–periodic Jacobi operators (for which $h(n)=0$ for $|n|\geq 2$), and then extend it to all $1$–periodic Hamiltonians by an approximation argument. The proof that the spectral flow equals $-{\mathcal{N}}(E)$ for Jacobi operator is postponed to the next Section. In this section, we explain how to generalize the result from Jacobi operators to any periodic Hamiltonian.
@@ A Jacobi operator is an operator $H$ acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ of the form
We will establish a version of Theorem 1.1 in the special case of Jacobi operators.
@@
We denote by $\widetilde{H}$ the corresponding operator acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{\ell N})$. Then $\sigma(H)=\sigma(\widetilde{H})$, and each Bloch band of $H$ becomes a set of $\ell$ bands for $\widetilde{H}$. In particular, ${\mathcal{N}}_{\widetilde{H}}(E)=\ell{\mathcal{N}}_{H}(E)$.
@@
$\widetilde{w}(x):=\begin{pmatrix}w(\ell x)&0&\cdots&0\\ 0&w(\ell x+1)&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&w(\ell x+\ell-1)\end{pmatrix},$
@@
${\mathrm{Sf}}\left(\widetilde{H}^{\sharp}(t),E,[0,1]\right)={\mathrm{Sf}}\left(H^{\sharp}(t),E,[0,\ell]\right)=\ell\,{\mathrm{Sf}}\left(H^{\sharp}(t),E,[0,1]\right).$
@@
${\mathrm{Sf}}\left(\widetilde{H}^{\sharp}(t),E,[0,1]\right)=-{\mathcal{N}}_{\widetilde{H}}(E)=-\ell{\mathcal{N}}_{H}(E),$
@@
This kernel has already been introduced in (8). Recall that $\|H-H_{\ell}\|_{\rm op}\to 0$ as $\ell\to\infty$ by Young’s inequality. Note also that $\|H^{\sharp}(t)-H_{\ell}^{\sharp}(t)\|_{\rm op}=\|H-H_{\ell}\|_{\rm op}$ is independent of $t$, hence goes to 00 uniformly in $t\in\mathbb{R}$. In particular, the spectrum of $H_{\ell}^{\sharp}(t)$ approaches uniformly the spectrum of $H^{\sharp}(t)$.
@@
${\mathrm{Sf}}\left(H^{\sharp}(t),E,[0,1]\right)=\lim_{\ell\to\infty}{\mathrm{Sf}}\left(H^{\sharp}_{\ell}(t),E,[0,1]\right)=-{\mathcal{N}}(E),$
@@
To start, recall all eigenvalues branches are $\nu$-Lipschitz. Unfortunately, we could not find a reference for this well-known result, so let us give a short proof of this fact, and prove the following: Let $t\mapsto A(t)$ be a continuous family of self-adjoint operators which is $\nu$-Lipschitz, in the sense $\|A(t)-A(s)\|_{\rm{op}}\leq\nu|t-x|$ (this implies that they all have the same domain), then all isolated branches of eigenvalues are also $\nu$-Lipschitz.
In the case where $A(t)$ is matrix valued (finite dimensional case), we first note that for all normalized $\Psi$, the map $t\mapsto\langle\Psi,A(t)\Psi\rangle$ is $\nu$-Lipschitz. Indeed, we have
$\left|\langle\Psi,A(t)\Psi\rangle-\langle\Psi,A(s)\Psi\rangle\right|=\left|\langle\Psi,A(t)-A(s),\Psi\rangle\right|\leq\|A(t)-A(s)\|_{\rm{op}}\|\Psi\|^{2}\leq\nu|t-s|.$
Since the $j$-th lowest eigenvalue $\lambda_{j}(t)$ is defined as a min–max combination of such $\nu$–Lipschitz functions, we conclude that all maps $t\mapsto\lambda_{j}(t)$ are $\nu$–Lipschitz.
We now turn to the general case. Let $t_{0}\in\mathbb{R}$, let $\lambda_{0}:=\lambda(t_{0})$ be an isolated eigenvalue of $A(t_{0})$, and let $m=\dim{\mathrm{Ker}}(A(t_{0})-\lambda_{0})$ be its multiplicity. By continuity of the spectrum [27], there is $\varepsilon>0$ and $\eta>0$ so that for all $t\in(t_{0}-\varepsilon,t_{0}+\varepsilon)$ the interval $(\lambda_{0}-\eta,\lambda_{0}+\eta)$ contains exactly $m$ eigenvalues of $A(t)$. We label the smallest eigenvalue as $\lambda_{1}(t)$ and $\lambda_{m}(t)$ for the largest one. In addition, by taking $\varepsilon$ and $\eta$ small enough, we may further assume that
$\begin{cases}{\rm dist}\left(\lambda_{0}-\eta,\sigma(A(t))\right)={\rm dist}\left(\lambda_{0}-\eta,\lambda_{1}(t)\right)=\lambda_{1}(t)-(\lambda_{0}-\eta),\\ {\rm dist}\left(\lambda_{0}+\eta,\sigma(A(t))\right)={\rm dist}\left(\lambda_{0}+\eta,\lambda_{m}(t)\right)=(\lambda_{0}+\eta)-\lambda_{m}(t).\end{cases}$
In particular, since $A(t)$ is self-adjoint, we have $\|\left(A(t)-(\lambda_{0}-\eta)\right)^{-1}\|_{\rm{op}}=\left(\lambda_{1}(t)-(\lambda_{0}+\eta)\right)^{-1}$. Let us prove that $t\mapsto\lambda_{1}(t)$ is $\nu$–Lipschitz on the interval $(t_{0}-\varepsilon,t_{0}+\varepsilon)$. We have
$\left|\lambda_{1}(t)-\lambda_{1}(s)\right|$ $=\left|\lambda_{1}(t)-(\lambda_{0}-\eta)\right|\cdot\left|\lambda_{1}(s)-(\lambda_{0}-\eta)\right|\left|\frac{1}{\lambda_{1}(s)-(\lambda_{0}-\eta)}-\frac{1}{\lambda_{1}(t)-(\lambda_{0}-\eta)}\right|,$
and the last term is also
$\left|\left\|\frac{1}{A(s)-(\lambda_{0}-\eta)}\right\|_{\rm{op}}-\left\|\frac{1}{A(t)-(\lambda_{0}-\eta)}\right\|_{\rm{op}}\right|\leq\left\|\frac{1}{A(s)-(\lambda_{0}-\eta)}-\frac{1}{A(t)-(\lambda_{0}-\eta)}\right\|_{\rm{op}}$
$\quad\leq\left\|\frac{1}{A(s)-(\lambda_{0}-\eta)}\right\|_{\rm{op}}\left\|\frac{1}{A(t)-(\lambda_{0}-\eta)}\right\|_{\rm{op}}\left\|A(t)-A(s)\right\|_{\rm{op}}$
We conclude that $\left|\lambda_{1}(t)-\lambda_{1}(s)\right|\leq\nu|t-s|$. The proof that $\lambda_{m}(s)$ is $\nu$–Lipschitz on the interval $(t_{0}-\varepsilon,t_{0}+\varepsilon)$ is similar. Finally, for any local continuous branch of eigenvalue $\lambda(\cdot)$ with $\lambda(t_{0})=\lambda_{0}$, we have
$\lambda_{1}(t)-\lambda_{0}\leq\lambda(t)-\lambda_{0}\leq\lambda_{m}(t)-\lambda_{0},\quad\text{so}\quad\left|\lambda(t)-\lambda(t_{0})\right|\leq\nu|t-t_{0}|.$
We therefore proved that for all $t_{0}$, there is $\varepsilon>0$ so that any continuous branch $\lambda(\cdot)$ of isolated eigenvalue satisfies $|\lambda(t)-\lambda(t_{0})|\leq\nu|t-t_{0}|$ for all $t\in(t_{0}-\varepsilon,t_{0}+\varepsilon)$. By compactness of $[t,s]$ for all $t,s\in\mathbb{R}$, we conclude that $\lambda(\cdot)$ is globally $\nu$–Lipschitz.
Now, let $E\in\mathbb{R}\setminus\sigma_{\rm{bulk}}$ with ${\mathcal{N}}(E)>0$. Since there is a nonzero spectral flow, $H^{\sharp}(t)$ has eigenvalues for at least some values of $t$. By shifting $t$ if necessary, we may assume $t=0$. We label this eigenvalue $\lambda_{0}(0)$, and denote by $\lambda_{0}(t)$ a corresponding continuous branch. We use this as a starting point to label the eigenvalue curves $\lambda_{k}(t)$ for $t\in[0,1]$ such that, for all $t\in[0,1]$
@@
and with the convention that $\lambda_{k}(t)$ equals the edge of the essential spectrum if there exists no $k$-th eigenvalue at this value of $t$. Since the spectrum is periodic, we have $\lambda_{k}(1)=\lambda_{j}(0)$ for some $j\in\mathbb{Z}$. The spectral flow equals $-{\mathcal{N}}(E)$, so there are exactly ${\mathcal{N}}(E)$ eigenvalues that cross $\lambda_{k}(1)$ upwards. We deduce that we actually have $\lambda_{k}(1)=\lambda_{k+{\mathcal{N}}(E)}(0)$ for all values of $k\in\mathbb{Z}$ such that $\lambda_{k}(0)$ is an eigenvalue of $H^{\sharp}(0)$. Then, the Lipschitz continuity of $t\mapsto\lambda_{k}(t)$ gives
$\lambda_{k+{\mathcal{N}}(E)}(0)-\lambda_{k}(0)=\lambda_{k}(1)-\lambda_{k}(0)\leq\nu.$
Hence, there are at least ${\mathcal{N}}(E)$ eigenvalues for $H^{\sharp}(0)$ in each interval of length $\nu$ in the gap. If $\nu$ is smaller than the width of the gap, the same argument shows that there is edge spectrum for any $t_{0}\in[0,1]$, and we can repeat the counting argument for this $t_{0}$.
@@ 2.4.2. Numerical simulations for the SSH chain
which corresponds to the evaluation of $w_{1}$ at the location of each of the two atoms in the first unit cell. Note that $w$ satisfy the assumption of a soft wall, and is $\nu$–Lipschitz.
@@
Then, for each eigenpair $(\lambda,u)$ with $\|u\|_{\ell^{2}_{L}}=1$, we check if the mass of $u$ is localized near the origin, or, more specifically, not localized at the boundary of our simulation box. This procedure discards the spurious edge modes due to the truncation. We detect that $u$ is a true edge mode whenever $\|u{\mathds{1}}(-L/2\leq n\leq L/2)\|_{\ell^{2}_{L}}\geq 3/4$.
@@
Figure 4. Numerics for the spectrum of $t\mapsto H^{\sharp}(t)$ in the SSH model with the potential $w_{1}(x)$ , for different values of $\nu$. From left to right, $\nu=0.5$, $\nu=1$, $\nu=5$ and $\nu=10$. The essential spectrum is displayed in grey, and the edge spectrum in red. For $E$ in the gap between the $2$ bands, ${\mathcal{N}}(E)=1$ and the spectral flow is $-1$, above the bands, ${\mathcal{N}}(E)=2$ and the spectral flow equals $-2$. When the Lipschitz constant increases, the curves become steeper.
@@ 3. Computation of spectral flows in the Jacobi case
Before we start, we emphasize that several techniques are available to study Jacobi operators, such as the ones involving transfer matrices. If $H$ is a Jacobi operator, then the equation $H\Psi=E\Psi$ implies that
@@
so that we can study the properties of $\Psi$ by studying the transfer matrix $T(E)$. However, this matrix is not always well-defined (in the previous formulation, we need $a$ to be invertible for instance). Actually, we were not able to recover our results using the formalism of transfer matrices, even in the case where $a$ is invertible. Our proof rather relies on a direct study of the full operator, and a variational argument.
@@ 3.1. Reduction to the cut case
$\forall x>0,\quad w_{\Sigma}(x)=0,\quad\text{while}\quad\forall x<-1,\quad w_{\Sigma}(x)\geq E+3C_{a,b}+1.$
@@
The difference of potentials $(w-w_{\Sigma})$ is continuous, vanishes for $x<x_{\Sigma}$, and goes to 00 as $x\to+\infty$. In particular, for all $t\in\mathbb{R}$, $(W-W_{\Sigma})(t)$ is compact, and the map $t\mapsto(W-W_{\Sigma})(t)$ is continuous and translation equivariant.
@@
$H^{\Sigma}(t)-H^{\sharp}(t)=W_{\Sigma}(t)-W(t),$
which is compact for all $t\in\mathbb{R}$. In addition, both operators $H^{\Sigma}$ and $H^{\sharp}$ are translation equivariant. By robustness of the spectral flow with respect to compact operators, see Lemma A.11, we deduce that
@@
Note that $K(1)=\tau_{1}K(0)\tau_{1}^{*}$, so the map $t\mapsto K(t)$ can be extended by translation equivariance, and this extension is continuous. For simplicity, in what follows, we will restrict the study to $t\in[0,1]$.
@@
$H_{\rm{cut}}^{\Sigma}(t):=H^{\Sigma}(t)-K(t),\quad\text{so that}\quad H^{\Sigma}(t)=H^{\Sigma}_{\rm{cut}}(t)+K(t).$
Since $t\mapsto H^{\Sigma}(t)$ and $t\mapsto K(t)$ are continuous and translation equivariant, so is $t\mapsto H_{\rm{cut}}^{\Sigma}(t)$. So, using again the robustness of the spectral flow with respect to compact, translation equivariant operators, see Lemma A.11, we find that
@@
These left and right operators are not periodic nor translation equivariant in $t$. Still, one can define the spectral flow as the net number of eigenvalues crossing the energy $E$ downwards. Since these operators are no longer translation equivariant, a priori this spectral flow could depend on the value of $E$ in the essential gap (see Appendix A). In any case, since $H_{\rm{cut}}^{\Sigma}(t)$ is block diagonal, we have
@@
$H^{\Sigma}_{R}(0)=\left(\begin{array}[]{| c c c c }\hline\cr b&a&&\\ a^{*}&b&a&\\ &a^{*}&b&\ddots\\ &&\ddots&\ddots\end{array}\right),\qquad\text{and}\qquad H^{\Sigma}_{R}(1)=\left(\begin{array}[]{| c c c c }\hline\cr\Sigma&0&&\\ 0&b&a&\\ &a^{*}&b&\ddots\\ &&\ddots&\ddots\end{array}\right).$
As noted before, the half-line operators are no longer translation equivariant: the presence of the boundary at 00 makes the model at $t=1$ not unitarily equivalent to the one at $t=0$. The key observation is that the relevant part of the spectrum is still periodic, since $H^{\Sigma}_{R}(1)$ is a direct sum of $\Sigma$ with (a shifted version of) $H^{\Sigma}_{R}(0)$. We will study such families of operators in the next section.